### v1

# Can Frontier LLMs Replace Annotators in Biomedical Text Mining? Analyzing Challenges and Exploring Solutions

Abstract
Large language models (LLMs) can perform various natural language processing (NLP) tasks through in-context learning without relying on supervised data. However, multiple previous studies have reported suboptimal performance of LLMs in biological text mining. By analyzing failure patterns in these evaluations, we identified three primary challenges for LLMs in biomedical corpora: (1) LLMs fail to learn implicit dataset-specific nuances from supervised data, (2) The common formatting requirements of discriminative tasks limit the reasoning capabilities of LLMs particularly for LLMs that lack test-time compute, and (3) LLMs struggle to adhere to annotation guidelines and match exact schemas, which hinders their ability to understand detailed annotation requirements which is essential in biomedical annotation workflow. To address these challenges, we experimented with prompt engineering techniques targeted to the above issues, and developed a pipeline that dynamically extracts instructions from annotation guidelines. Our findings show that frontier LLMs can approach or surpass the performance of state-of-the-art (SOTA) BERT-based models with minimal reliance on manually annotated data and without fine-tuning. Furthermore, we performed model distillation on a closed-source LLM, demonstrating that a BERT model trained exclusively on synthetic data annotated by LLMs can also achieve a practical performance. Based on these results, we explored the feasibility of partially replacing manual annotation with LLMs in production scenarios for biomedical text mining.
1 Introduction
1.1 Background: Annotation Paradigms in Biomedical Text Mining
Scientific literature contains a wealth of valuable information, presenting both opportunities and challenges for biological and medical research. Extracting knowledge from vast amounts of unstructured biomedical literature is critical for building biological databases, enabling researchers to uncover new insights and conduct data-driven studies. For example, the Gene Ontology (GO) project builds a comprehensive database by extracting information about genes and protein functions from scientific literature through a combination of manual curation and automated tools[1, 2]. This ontology database provides detailed classifications of biological processes, molecular functions, and cellular components of genes. Meanwhile, the ever-growing volume of published biomedical literature has exacerbated the issue of information overload, making it increasingly difficult for scientists, medical professionals, and the general public to quickly and accurately retrieve required information. For example, since the onset of the COVID-19 pandemic in 2020, over 400,000 papers related to COVID-19 and SARS-CoV-2 have been published in at least 8000 academic journals[3]. This urgents the need for effective automated text mining tools.
Extracting information from biological literature relies heavily on annotation conducted by human experts. Typically, curators define the annotation tasks and objectives, then collaborate with domain experts to develop annotation guidelines. Subsequently, human annotators conduct the annotation based on these guidelines[4, 5]. In many production use cases, manual annotation remains essential for continuously updating data to maintain annotation accuracy. Examples include projects like UniProtKB/Swiss-Prot and the Gene Ontology[6]. The labor-intensive nature of manual annotation has driven the adoption of natural language processing (NLP) methods in biomedical text mining[7]. These methods enable the automated extraction of valuable information from large volumes of unstructured text, including tasks such as identifying key entities (Named Entity Recognition, NER) and uncovering relationships between them (Relation Extraction, RE). Additionally, text classification plays a critical role in categorizing and organizing this information effectively. A well-known example of automated text mining in biological domain is PubTator, which applies machine learning techniques to annotate entities and relationships in biological literature, thereby assisting researchers in efficiently accessing and interpreting information[8, 9]. Another example is LitCovid, a web system for tracking the latest publications and research on COVID-19, which uses machine learning to classify newly published documents automatically[10, 11, 12]. However, even in such automated annotation workflows, human annotator are also required for building annotated datasets used for supervised learning models.
In the broader field of biomedical natural language processing (BioNLP), generative tasks such as medical question answering and text summarization also play a notable role. Nevertheless, the present study primarily focuses on information extraction, a category often referred to as discriminative tasks. Our focus on these tasks stems from the observation that large language models continue to face challenges in this domain, as will be discussed in detail in Section 1.2
Figure 1: The Reliance of Biomedical Text Mining Tasks on Human Annotators
a. This panel illustrates two paradigms of biomedical text mining in real-world applications: (i) Continuous annotation by human experts, commonly seen in scenarios requiring high annotation precision, such as database curation. (ii) The constructions of supervised datasets by human experts, which are then used to train machine learning models for annotating large-scale or continuously produced texts.
b. This panel demonstrates the dependency of machine learning performance on the amount of training data in the multi-label text classification dataset LitCovid. Achieving satisfactory model performance requires the manual annotation of hundreds or even thousands of data samples.
1.2 Background: Studies on Language Models in BioNLP
Early biomedical text mining relied primarily on rule-based methods, statistical techniques, and early machine learning approaches[13, 14, 15, 16, 16]. In recent years, language models based on neural networks have achieved dominant performance: BERT based on Transformer encoder architecture have shown significant advances by leveraging pre-training on unannotated corpora to understand natural language and contextual knowledge[17, 18, 19]. These models can then be fine-tuned on smaller supervised datasets to perform various specific biomedical tasks. Despite their strong performance across diverse biomedical text mining applications, BERT-like models face a critical limitation that they depend on supervised data. As shown in Figure 1 b., the size of the corpus and the availability of labeled data directly influence the performance of these models[20].
The emergence of large language models (LLMs), exemplified by ChatGPT, seems to offer a potential solution to this challenge. LLMs are typically based on Transformer decoder architecture and are pre-trained on large-scale copora using next-token prediction to develop text generation capabilities[21][22][23]. Additionally, LLMs undergo alignment training like instruction-tuning to improve their ability to follow human instructions, engage in logical reasoning, and adapt to diverse tasks without requiring supervised data or specialized training[24].
As early as 2022, studies began applying models like GPT-3 to information extraction, showing promising results in clinical domain[25]. Following the release of ChatGPT, Qijie Chen et al. benchmarked GPT-3.5 on various tasks in the biomedical literature[26]. While the model outperformed baselines in Q&A tasks, its performance on information extraction tasks lagged significantly behind those of fine-tuned models. Similarly, a comprehensive evaluation of LLMs across 26 datasets conducted by Jahan et al. confirmed that these models struggle to match the performance of fine-tuned BERTs in discriminative tasks[27]. Further studies, such as those by Qingyu Chen et al. highlighted that even advanced proprietary models like GPT-4 which excel in biomedical generative tasks, exhibit notable performance deficits compared to fine-tuned BERT models in discriminative tasks[28]. In some datasets, GPT-4 achieved only 40% of the SOTA performance (e.g., 0.33 compared to 0.79). They furthermore analyzed the errors made by LLMs, but those analysis primarily focused on output format issues, without delving into the reasoning processes of the models.
1.3 Contribution of Our Work
We focus on three representative tasks in biomedical text mining—named entity recognition (NER), relation extraction (RE), and multi-label text classification. We first reproduced the previous evaluations of LLM in biological domain to analyze the failure patterns, then introduced series of prompt engineering methods and a information retrieval pipeline specially targeted those issues we identified:
(1) We revealed a significant discrepancy between the distribution of LLM outputs and the ground truth label across various dimensions such as the number and distribution of output labels, average span length of entity mentions. In contrast, models fine-tuned through supervised learning exhibit much smaller discrepancies. We assume this difference arises because supervised training implicitly captures dataset-specific features. To address this issue, we introduced an automatic prompt optimization techinique targeted to narrowing the gap between LLM-predicted labels and the correct labels.
(2) Structured batch outputs are often favored for their convenience in parsing and downstream analysis in biomedical text mining tasks. However, this approach may unintentionally limit the ability of LLMs to showcase their reasoning processes during inference, which, as observed in our analysis and other studies, can negatively affect their overall performance[29, 30]. To mitigate that, we developed a two-step inference approach targeted to preserve chain-of-thought reasoning during structured outputs[31]. Additionally, in the later stages of our study, several commercial models emerged that support test-time computation, enabling automated reasoning before the final output. We also conducted evaluations on one of the representative reasoning models.
(3) Our analysis of LLM outputs highlights that many failures stem from a lack of understanding of task requirements. Annotation guidelines, which are often several pages to tens of pages long, specify crucial aspects such as task requirements, annotation categories, hierarchies, and granularities. They also provide examples, counterexamples, and instructions for handling edge cases[4, 5]. These details that are critical for accurate annotation but were largely overlooked in prior evaluations of LLMs. Thus, we explored incorporating annotation guidelines into prompts to address the issue of LLMs lacking an inherent understanding of task-specific requirements. Given that annotation guidelines are often lengthy, simply appending them to the prompt resulted in suboptimal performance, as LLM struggles to identify the most relevant sections. Baed on that, we manually divided the annotation guidelines into smaller text chunks and developed a Retrieval-Augmented Pipeline to dynamically retrieve the most relevant sections based on the task context.
Utilizing the above approach, we achieved significant improvements in Biological Text Mining of LLM performance, surpassing the SOTA on 2 datasets and demonstrating practical performance levels on the remaining.
Though some studies have shown that fine-tuning (including instruction tuning) may also be an effective approach[32, 33], the primary reason we did not adopt a fine-tuning approach is the substantial computational resources required for fine-tuning and deploying a large language model. Moreover, Nori et al. demonstrated that advanced prompt engineering can significantly enhance the performance of LLM and enabled a general-domain LLM to outperform domain-specific models[34]. Some reviews have also highlighted similar trends in related studies[35, 36, 37]. Additionally, inference-side scaling laws have garnered significant attention recently[38, 39, 40]. For example, GPT-o1 has shown exceptional performance in complex reasoning tasks by automatically generating reasoning steps before producing outputs[41].
In real applications, deploying a LLM presents challenges. LLM demands significant GPU memory, and the attention mechanism used in GPT architectures has a quadratic order of memory and computation complexity with respect to the sequence length, which results in high inference costs[17, 21]. Cloud-based solutions, while feasible, incur considerable financial costs and pose potential data security risks. To mitigate these challenges, we adopted a workflow inspired by production use scenarios: human-annotated data is used to train machine learning models, which then continuously perform inference. We treated the LLM together with its output parser as a discriminative model and distilled it into a lighter model like BERT by a style of weak supervision learning. For the HoC and BC5CDR-RE datasets (which are a multi-label text classification dataset and a relation extraction dataset, respectively), we collected literature published after April 2024 following the same procedure as in the original studies[42, 43]. We then used the LLM to generate annotated data, which were used for fine-tuning a BERT model. The distilled BERT model achieved performance close to the previous state-of-the-art (SOTA).
Our results suggest that for emerging text mining demands, relying solely on LLM inference in production environments or using an LLM+BERT distillation pipeline may be a pratical replacement of the widely adopted workflow of human annotation followed by machine learning.
2 Problem Analysis: LLM in Biomedical Text Mining
2.1 Biomedical Tasks and Datasets
We focus on three key tasks: named entity recognition (NER), relation extraction (RE), and multi-label text classification (CLS). Biomedical named entity recognition aims to extract biomedical named entities such as genes, sequences, drugs, and treatments from text. For example, "Mutational analysis of CYP2C8 in hypertension using a new technology." In this case, we would identify: "CYP2C8" as "Gene", "hypertensive" as "Disease". Biomedical relation extraction involves identifying relationships between biomedical entities, such as drug-drug interactions. This task typically requires extracting (Entity 1, Entity 2, Relationship Type) triples. For example, in "The combination of warfarin and aspirin increases the risk of bleeding.", expected outputs include (Warfarin, Aspirin, Adverse Effect). Biomedical multi-label text classification involves assigning a document (sentence or article) to one or more labels. For instance, the possible labels of "Metformin is effective in controlling blood glucose levels and is also associated with gastrointestinal side effects." could be Effectiveness, Adverse Effect.
Table 1 summarizes the datasets used in our experiments along with their various characteristics. For our baselines, we selected well-known SOTA foundation models such as BioBERT[44] and BioLinkBERT[45], which have demonstrated strong performance across a wide range of tasks. However, since we were unable to identify any prior study about BC5CDR-RE dataset utilizing BERT-based models, we conducted our own fine-tuning and evaluation of BioBERT on this dataset in addition to reporting the results of the current SOTA result which was achieved by a fine-tuned BioGPT[46].
We adopt evaluation metrics that were commonly adopted in prior research. For multi-label text classification and relation extraction, evaluation metrics include Macro Precision, Macro Recall, and Macro F1-score. For named entity recognition task, the evaluation is conducted using entity-level metrics. Their definitions are as follows:
  • Macro-Level Metrics: Macro Precision, Macro Recall, and Macro F1-score provide a class-agnostic evaluation of model performance by averaging the metrics across all classes without weighting by the number of instances per class. This ensures that all classes, regardless of size, contribute equally to the evaluation. Let $C$ denote the total number of classes:
    $\text{Precision}=\frac{1}{C}\sum_{i=1}^{C}\frac{\text{True Positives}_{i}}{\text{True Positives}_{i}+\text{False Positives}_{i}}$
    $\text{Recall}=\frac{1}{C}\sum_{i=1}^{C}\frac{\text{True Positives}_{i}}{\text{True Positives}_{i}+\text{False Negatives}_{i}}$
[…]
  • Entity-Level Metrics: Metrics are computed by aggregating predictions and ground truths for all entities across the entire dataset. Let CorrectEntities be the number of correctly predicted named entities, PredictedEntities be the total number of predicted entities, and GoldEntities be the total number of gold-standard (actual) entities. Then:
    $\text{Precision}=\frac{\text{CorrectEntities}}{\text{PredictedEntities}},\quad\text{Recall}=\frac{\text{CorrectEntities}}{\text{GoldEntities}},\quad\text{F1}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
Table 1: Datasets and Their Key Features
Tasks Dataset
Datasize & Split
(Train/Valid/Test)
Evaluation
Metrics
Baselines
Annotation
Guideline
Named Entity Recoginization
BC5CDR-Chemical,
BC5CDR-Disease[43]
4560/4581/4797 Entity-Level BioLinkBERT-Large Y (8 Pages)
NCBI-Disease[47] 5424/923/940 BioBERT-Large v1.2 Y (5 Pages)
Relation Extraction BioRED[48] 4873/1162/1163 Macro-Level BiomedBERT Y (16 Pages)
BC5CDR-RE[43] 500/500/500 BioLinkBERT-Large, BioGPT Y (8 Pages)
Multi-label Classification LitCovid 24960/6239/2500 Macro-Level BioBERT-Large v1.2 N
HoC 1108/157/315 BioLinkBERT-Large N
  • When fine-tuning the baseline model BioBERT on the BC5CDR (RE) dataset, our initial exploration yielded catastrophic results as the model overwhelmingly predicted negative labels. To address this, we randomly sampled and reduced the number of negative labels in the training set. We achieved the results presented in the table when the ratio of negative to positive labels was reduced to 2:1,
  • Randomly sampled 500 test samples for LLM evaluation due to API cost constraints.
2.2 Challenges Faced by LLM
Figure 2: Models fine-tuned via supervised learning exhibit label distributions that align more closely with the training set and ground truth.
a. In the BC5CDR (RE) dataset, the training and test sets have an average of 1.96 and 2.16 CDI relationships per test instance, respectively. The predictions of a supervised BERT model average 2.27 CDI relationships per test instance, aligning more closely with the dataset’s characteristics. In contrast, GPT-4 predicts an average of 2.88 relationships, despite achieving higher performance (Macro-F1: 0.69 vs. 0.44).
b. Label frequency distribution in the LitCovid multi-label text classification dataset.
c. Average word count per entity mention in the BC5CDR-Disease named entity recognition dataset. GPT-4 outputs tend to select fewer words per entity, reflecting possible ambiguity in defining annotation boundaries.
d. Frequency distribution of relationship types in the BioRED relation extraction dataset.
In prior studies evaluating LLMs for biomedical text mining, the prompts typically included brief descriptions of the task type, labels, and output format. Building on these works, we reproduced their approaches using GPT-4 (API version: 0613[49]), consistent with Qingyu Chen’s research[28]) under a random 3-shot setting. The detailed results are presented in Table 2. Our evaluation scores are comparable to or slightly higher than, previous studies, which may be attributed to subtle differences in prompt text and improved robustness of the output parsers. For datasets that were previously only evaluated with GPT-3.5 or other LLMs, we observed improved results, likely due to the superior performance of GPT-4 itself[26, 27]. However, GPT-4 still significantly underperforms BERT-based models across all datasets except BC5CDR-RE.
To investigate the characteristics of outputs from LLMs, we analyzed the following statistics:
  • For named entity recolonization, we examined the number of entities per test instance and the average word length of each entity across the whole test set.
  • For relation extraction, we evaluated the number of relationships per test instance and the distribution of relationship categories.
  • For multi-label text classification, we analyzed the number of labels per test instance and the distribution of each label.
We selected four representative results to display in Figure 2, while additional detailed results are provided in Figures S1. From these results, we observe that across various feature metrics, BERT-based models generally align more closely with the ground truth labels in both the training and test sets compared to the outputs of GPT-4. Notably, GPT-4 achieved performance comparable to the fine-tuned state-of-the-art model (BioGPT) and outperformed fine-tuned BioBERT on the BC5CDR-RE dataset. However, the distribution of relations extracted by GPT-4 remained further from the test set distribution (one-sided Mann-Whitney U test, p-value = 9.25e-05, Wasserstein distance = 0.356) compared to BioBERT, which exhibited a closer match to the test set distribution (Wasserstein distance = 0.180).
Considering the reported superior performance of advanced closed-source models like GPT-4 in other generative tasks in biomedical domain and reports suggesting that their training data likely includes extensive biomedical literature[50, 51, 52, 53, 54], it is plausible that these models possess robust capabilities in understanding biomedical documents. This includes a strong grasp of domain-specific terminology, ambiguous expressions, and biomedical abbreviations. Given this, we hypothesize that the shortcomings of large language models (LLMs) in biomedical text mining tasks may partly stem from insufficient understanding of the datasets themselves. To validate this hypothesis, we conducted a case study focusing on the failures of LLMs in these tasks. These failures can be categorized as follows: 1. Errors that show some degree of reasonableness but fail to meet the detailed requirements about annotation categories, hierarchies, and granularities, or represent exceptions explicitly noted in the guidelines. 2. Errors arising from insufficient semantic understanding. 3. Parsing errors.
Table 2: Major Failure Patterns and Case Studies
Failure Pattern Explanations and Examples
1. Semantically Plausible but Inconsistent with Annotation Guidelines In many instances, predictions generated by GPT-4 exhibit a degree of rationality and may initially appear accurate. However, these predictions fail to conform to the specific and detailed requirements stipulated by the annotation guidelines, resulting in misalignment with the true labels. This category of errors can be further subdivided into three subtypes, though the boundaries between these subtypes are blurred: (a) Incorrect Annotation Hierarchy.
In the NCBI-Disease dataset, both mentions of specific diseases and disease classes (families of specific diseases) should be annotated. However, GPT-4 using prompt same as prior research often neglected disease class entities. For instance, in the sentence: “The RB1 gene mutation was investigated in a child with ectopic intracranial retinoblastoma using DNA obtained from both the pineal and retinal tumours of the patient.”(PMID: 9400934[55]) the LLM properly annotated all other disease mentions but failed to label “pineal and retinal tumours” as a disease class. In the BioRED (RE) dataset, only specific relationships (e.g., disease–chemical) among five defined entity types should be annotated. Nevertheless, the LLM labeled “dexmedetomidine” and “patient” as having an “Association” relationship in: “A comparison of severe hemodynamic disturbances between dexmedetomidine and propofol for sedation in neurocritical care patients.”(PMID: 24717468[56]) Where according to the guidelines, “patient” should be marked as a “Species” entity, which should not be linked in any relationship with a chemical entity in this dataset context.
(b) Incorrect Annotation Category.
In BioRED (RE) dataset, relationships must be classified into categories like “positive correlation,” “negative correlation,” or “association.” However, the LLM sometimes misinterpreted the direction of effect. Consider this sentence: “…intravenous lidocaine infusion or a high dose of mexiletine …controlled the ventricular tachycardia.”(PMID: 15485686[57]) The LLM labeled the pair (mexiletine, ventricular tachycardia) as “positive correlation.” While it is true that mexiletine has a beneficial therapeutic effect on this disease, the guidelines specify that when a drug is used to treat a disease or reduces its severity, it should be labeled as “negative correlation”.
(c) Incorrect Boundary or Granularity.
Certain datasets (e.g., BC5CDR-Disease) require separate mention annotations for a full disease name and any parenthetical abbreviations. However, The LLM occasionally merges both into a single entity mention. For instance, from “Parkinson’s disease (PD),” two separate mentions (“Parkinson’s disease” and “PD”) are required, but the LLM collapses them into a single mention.
(d) Neglecting Special Exceptions.
According to the NCBI-Disease guidelines, words like “disease”, “syndrome”, “deficiency”, or “complication” alone should not be annotated. However, in the following text: “Low levels of beta hexosaminidase A in healthy individuals with apparent deficiency of this enzyme.”(PMID: 941901[58]) the LLM incorrectly annotated “deficiency” as a disease despite the guideline explicitly prohibiting such labeling. For the following sentence in BioRED (RE) dataset: “The acute toxicity of OPs is the result of their irreversible binding with AChEs in …”(PMID: 15036754[59]) LLM annotated “OP” and “toxicity” as having a “negative correlation,” even though the annotation guideline requires not to annotate the chemicals with the disease concept-“toxicity” directly, but to annotate the disease due to the “toxicity” and “drug-induced” instead.
2. Errors due to Inadequate Semantic Comprehension These are pure reasoning mistakes by the LLM and are not necessarily tied to annotation guidelines. For instance, in the following article snippet from BC5CDR (RE) dataset: “Late-onset scleroderma renal crisis induced by tacrolimus and prednisolone: a case report. Scleroderma renal crisis (SRC) is a rare complication of systemic sclerosis (SSc) but can be severe enough to require temporary or permanent renal replacement therapy. Moderate to high dose corticosteroid use is recognized as a major risk factor for SRC. Furthermore, there have been reports of thrombotic microangiopathy precipitated by cyclosporine in patients with SSc.”(PMID:22836123[60]) The correct chemical–disease interactions (CDI) that should be extracted include: Cyclosporine–microangiopathy, Corticosteroid–SRC/SSc, tacrolimus–SRC/SSc. However, the LLM omitted (tacrolimus, SRC/SSc) and misidentified the disease associated with “cyclosporine” as SRC/SSc instead of microangiopathy. Interestingly, in a simpler prompt setting (e.g., “retrieve all the CID pairs from the text”), the LLM produced the correct pairs in a more free-text style. This suggests that forcing the LLM to output fully structured annotations may disrupt its internal reasoning processes. This observation motivated the technique we propose in Section 3.1.3.
3. Parsing Errors A parsing error occurs when a large language model (LLM) fails to produce output in the specified structured format, rendering the output parser unable to process the result correctly. While models like LLaMA2-13B have frequently reported these inconsistencies according to Q. Chen et al[28, 61], we have not encountered such errors outside the development phase. This can be attributed to that models like GPT-4 benefit from extensive training focused on generating structured outputs, as well as that we designed a robust output parser capable of effectively capturing edge cases.
3 Methodology
To address the challenges identified in Section 2.2, we applied a targeted prompt engineering strategy for biomedical text mining tasks. This strategy includes general techniques like dynamic few-shot prompting and some targeted methods like two-step inference for maintaining both reasoning processes alongside structured output, automatic prompt optimization to account for distributional differences between LLM outputs and ground truth labels, and a workflow for dynamically extracting annotation guidelines to provide detailed instructions to LLMs.
The experiments mainly utilized the GPT-4o model (2024-08-06), which was one of the most advanced and widely used models. Additionally, in the later stages of this study, several LLMs supporting test-time compute were released, which are not constrained by missing reasoning steps due to formatting limitations. Based on that, we also evaluated OpenAI o1-mini which is a representative model in this category (4.2).
To ensure that the observed performance improvements were not solely due to advancements in the underlying LLM itself, we also conducted experiments on the LitCovid and BC5CDR-Chemical datasets using historical model consistent with previous studies (see Appendix 7.2). Moreover, we applied model distillation techniques to the HoC and BC5CDR-RE tasks. This involved annotating papers published in April 2024 using LLMs to create augmented datasets, which were subsequently used for fine-tuning BERT-based models.
In the following, we describe these components in detail.
3.1 Prompt Engineering
3.1.1 Dynamic Few-shot Prompting
In-Context Learning (ICL) is a foundational capability of large language models (LLMs), enabling them to solve new tasks by providing instructions or a few examples of the task in the input[22]. Research has shown that the choice of examples used in few-shot prompting significantly affects model performance[50]. Based on that, We implemented a semantic-based dynamic few-shot prompting approach, which is a widely adopted technique across diverse tasks.
As showned in Figure 3 , we embedded each document (a combination of its title and abstract) in the dataset to a high-dimensional embedding and stored these embeddings in a vector database. In the inference phase, the $K$ nearest documents in the embedding space, based on cosine similarity, were retrieved as examples along with their expected outputs. We used OpenAI’s text-embedding-3-small model as the embedding model, Chroma for the vector database, and set $K=5$ for our experiments.
Figure 3: Workflow of the Dynamic Few-shot Prompting
3.1.2 Strategies for Structured Output
In our study on biomedical multi-label text classification, we adopt JSON as the standard output format. JSON is widely recognized as a structured data format for output generation in large language models (LLMs) and is explicitly supported by APIs such as OpenAI’s. In our implementation for text classification tasks, each key in the JSON structure corresponds to a label, and its associated value indicates whether the label is selected.
However, JSON is not well-suited for Named Entity Recognition (NER) and Relation Extraction (RE) tasks. For NER, the conventional approach involves the BIO format[62], which uses the labels B (Beginning), I (Inside), and O (Outside) to denote whether a token—typically a word or subword—marks the start of an entity, is inside an entity span, or lies outside any entity span. This approach aligns with the pre-LLM paradigm for NER, where the task is framed as sequence labeling: predicting the properties of each individual token. Prior studies such as those by I. Jahan et al., which evaluated LLMs in biomedical NER tasks using the BIO format, showed a very limited performance[27]. LLMs, by design, excel at generating coherent text rather than making predictions for individual tokens. To address this, S. Wang et al. developed a method that uses delimiters to enclose entities, achieving promising results[63]. Building on this insight and considering that contemporary LLM training datasets often include a significant amount of HTML-formatted text, we adopted an HTML tagging scheme for NER tasks. In our approach, entities are marked using tags such as: <Type = "$\{entity\ type\}$"> $\{entity\ mention\}$ </Type>
For RE tasks, which typically involve two entities and their relationship, we represent relationships using tuples. In the BC5CDR-RE task, where only one relationship type (Chemical-Disease Interaction, CDI) is involved, we use a format like: ($entity_{1}$, $entity_{2}$). In contrast, for BioRED, which includes multiple relationship types, we use the format of triples: ( $entity_{1}$, $entity_{2}$, $relation\_type$ ).
3.1.3 Two-step Inference to Maintain Both Reasoning and Structured Outputs
Chain-of-Thought (CoT) is a prompting technique designed to enhance reasoning by incorporating intermediate steps before arriving at the final answer [31]. By breaking down complex problems into smaller, more manageable steps, CoT has been shown to improve the accuracy of foundational model outputs. Frontier large language models (LLMs) are often trained on datasets that include reasoning processes, enabling them to autonomously decompose problems and engage in step-by-step reasoning [64]. As a result, this capability may naturally emerge even in the absence of explicit instructions for problem decomposition in the prompt. However, in biomedical text mining tasks, models are typically required to extract structured results directly from query text in bulk, rather than generating unstructured, difficult-to-parse textual responses.
Some recently released models that support test-time computation do not inherently suffer from this issue. However, for widely used models from the pre-test-time computation era, which offer faster inference (such as GPT-4o), structured constraints often force the model to bypass intermediate reasoning processes. Such omissions can degrade model performance, as demonstrated in Table 2 under “2. Errors due to Inadequate Semantic Comprehension.” Quantitative analyses by Z. Tam et al. further corroborate these findings, highlighting the impact of insufficient reasoning on task outcomes [29].
To address this challenge, we introduce a two-step approach that allows LLMs to first generate complete reasoning and thought processes, and subsequently extract structured answers from them (Figure 3.1.3). For multi-label biomedical text classification tasks requiring JSON-formatted outputs, we inserted an "intermediate steps" schema before the final output in OpenAI’s structured output[65]. For other tasks such as NER and RE which demand alternative output formats, and for the comparison test in LitCovid dataset using the legacy GPT-4 model where the strcutured ouput was not supported in the API, we implemented this functionality independently.
captionWorkflow of the Two-step Inference to Maintain Both Reasoning Step and Structured Outputs
3.1.4 Automatic Prompt Optimization
In Section 2.2, our analysis reveals a discrepancy between the distribution of labels predicted by LLM and the ground truth labels, and the inability of capturing such implicit features may contribute to their suboptimal performance. To address this issue, we aim to optimize prompts by narrowing the gap between predicted and correct label distributions.
Numerous automatic prompt engineering methods have been developed[37, 36], and we adopt and modify R. Pryzant’s pseudo-“gradient descent” approach which was inspired by gradient descent but the target of optimizatation is the prompt instead of the model parameters[66]. In this method, we identify distributional gap between LLM predictions and ground truth distributions on small batches of data, along with errors in the LLM’s output, to construct a natural language "gradient" that critiques the current prompt. By editing the prompt in the semantic direction opposite to this gradient, we aim to propagate the gradient into the prompt and identify a prompt that optimizes the desired metric (Macro-F1 score) at the dataset level. During the prompt refinement process, multiple candidate prompts are generated, and beam search is employed as the search algorithm. In each iteration, the top k prompts with the highest scores are retained as candidates for the next round, while lower-scoring prompts are discarded.
However, this method is computationally expensive, primarily due to the need for iterative batch inferences. To achieve reliable estimates of distribution differences, we set the batch size to 100. Additionally, implementing this approach for different tasks requires significant coding effort. As a result, our experiments were limited to the LitCovid and BC5CDR-Chemical dataset. For each iteration, we generated three pseudo-gradients along with corresponding candidate prompts during the "propagation" step. In the beam search, the top three candidates were retained in each iteration, and a total of 10 iterations were performed.
Figure 4: Automatic Prompt Optimization Using Pseudo "Gradient Descent"
3.2 Instruction Retrieval from Annotation Guideline
Our analysis in Section 2.2 reveals that limitations of LLMs in biomedical text mining arise primarily from their inability to fully comprehend the dataset-specific detailed annotation requirements. At the early stage in our research, we attempted to integrate complete annotation guidelines into the prompts for the NCBI-Disease (NER) and BioRED (RE) datasets to address this issue. However, this approach yielded negligible performance improvements. A plausible explanation is the excessive length of these guidelines, which makes it difficult for LLM to extract the necessary information within such a long context.
Based on that, we explored dynamically extracting relevant entries from the annotation guidelines for each query case. A common approach for such tasks is Retrieval-Augmented Generation (RAG), which integrates retrieval mechanisms with external knowledge bases to provide LLMs with accurate contextual information[67, 68]. While RAG techniques are already implemented in advanced commercial generative models, we could not find a readily available framework suitable for our specific needs. Therefore, we developed a custom rule-based pipeline tailored to these requirements.
For example, in the BioRED (RE) dataset, the annotation guidelines span 16 pages, with around 10 pages dedicated to relationship extraction. Within these, the “relation pairs” section accounts for about 4 pages, and the “relation types” section covers 6 pages. These requirements are further divided into seven subsections based on entity types (e.g., “Disease-chemical,” “Disease-variant & Disease-gene”). Specific rules for edge cases are either grouped in a separate “Others” subsection or scattered across the other subsections. To streamline the process, we segmented and reorganized the guidelines into structured text chunks, as shown in Figure 5 "Knowledge Construction Phase".
During inference, we generate all possible combinations of entities from the entity list and group them based on their types. We process all pairs within the same type in a single query instead of querying each entity pair individually to reduces API costs. Each test query is carried out in three steps (as illustrated in Figure 5 "Inference Phase").
  1. 1.
    Relation(Entity Pair) Prediction: We first predict whether a relationship exists between two entities. All candidate entity pairs within the same type are provided as input, and we incorporate the corresponding Relation Pair Requirements into the prompt to instruct the model.
  2. 2.
    Relation Type Classification: We then predict the relationship type for the identified entity pairs. Here, the Relation Type Requirements are appended to the prompt to refine the prediction.
  3. 3.
    Rule Validation: In the final step, we perform rule validation to verify if the predicted entity pairs and their relationship types require special handling. This step ensures the outputs align with the annotation guidelines and satisfy the necessary constraints.
For the BC5CDR-Chemical/Disease (NER) and NCBI-Disease Named Entity Recognition (NER) datasets, we adopt a similar approach. First, we provide LLM with requirements for identifying and labeling entity mentions to guide the model annotates the mentions accordingly. Then, we integrate instructions for handling edge cases to the prompt and guide LLM to check if the annotations adhere to the specified requirements. The BC5CDR-Chemical/Disease dataset includes two types of entities—Chemical and Disease—which required separate executions of the annotation steps. For the NCBI-Disease dataset, which contains only Disease entities, the steps were executed once. For the BC5CDR (RE) dataset, we simplified the process because the annotation guideline spans less than half a page. We embedded the guideline directly into the prompt, and let the LLM to produce results in a single query.
Figure 5: Instruction Retrival Workflow in BioRED (RE) Dataset
3.3 Model Distillation
To explore the potential of large language models (LLMs) in replacing human annotators in the widely adopted “manual annotation–model training–inference” workflow prevalent in real-world applications, we treat the LLM combined with its output parser as a discriminative model. We then conduct a model distillation process from GPT-4o to a BERT-based model. This process leverages data augmentation and weak supervision, which are fundamental strategies in LLM’s model distillation[69].
Specifically, we utilized an LLM to generate labeled data for unannotated research articles, aiming to train a BERT model for the Hallmarks of Cancer (HoC) dataset. We utilized the same search keyword set as the original HoC paper and retrieved PubMed articles published after April 2024 via the NCBI’s API (E-utilities)[70, 71]. This specific timeframe was chosen to minimize the risk of overlap between the retrieved articles and the LLM’s training data. Furthermore, this approach mimics a real-world scenario where novel annotation needs arise, requiring the labeling of newly published articles. After doing the annotation utilizing GPT-4o with two-step inference, we then filtered the retrieved articles to exclude those with no positive labels during the multi-label classification process, and constructed a LLM-annotated corpus of 3,000 articles. This dataset was then used to fine-tune a BioLinkBERT-Large model.
Similarly, for the BC5CDR-RE dataset, we followed the same procedure as in the original study that for each articles in BC5CDR corpus, we retrive all the related articles published after April 2024 by E-utilities. Using the titles and abstracts of these papers, we first performed named entity recognition (NER) to extract entity lists, followed by relation extraction (RE) to identify chemical-disease interactions (CDI) utilizing GPT-4o with two-step inference + instruction retrieval from the annotation guideline. From this set, we built an LLM-augmented corpus of 2361 articles including 5930 CDI records. This corpus was then used to fine-tune the BioLinkBERT-Large model. Negative samples were created by pairing Chemical and Disease entities that did not exhibit CDI relationships, and the number of negative samples was reduced through random sampling to maintain a 2:1 ratio of negative to positive samples.
4 Results and Discussion
4.1 Evaluation of GPT-4o Performance With Proposed Methods
Tables 3, 4, and 5 summarize the results of our experiments for GPT-4o across multi-label text classification, named entity recognition, and relation extraction tasks with targeted prompt engineering. Information about baseline models can be referred in Table 1. Additionally, five out of six failure cases listed in Table LABEL:tab:failures (PMIDs: 24717468, 15485686, 941901, 15036754, 22836123) were successfully addressed using the “Two-step Inference + Instruction Retrieval from Annotation Guideline” approach. The remaining case (PMID: 9400934) was resolved using the “Dynamic 3-shot + Instruction Retrieval from Annotation Guideline” method.
Table 3: Results on Multi-label Text Classification Tasks
LitCovid HoC
Precision Recall F1 Precision Recall F1
Basic Instruction 0.663 0.892 0.744 0.705 0.896 0.797
Dynamic 3-shot 0.739 0.860 0.798 0.720 0.857 0.786
Dynamic 3-shot + Automatic Prompt Optimization 0.815 0.839 0.822 - - -
Two-step Inference 0.831 0.909 0.873 0.827 0.865 0.841
Baseline (SOTA Foundation model) 0.870 0.849 0.861 0.887 0.864 0.873
Table 4: Results on Named Entity Recoginization Tasks
BC5CDR-Chemical BC5CDR-Disease NCBI-Disease
Precision Recall F1 Precision Recall F1 Precision Recall F1
Basic Instruction 0.814 0.739 0.775 0.629 0.441 0.519 0.710 0.645 0.676
Dynamic 3-shot 0.828 0.789 0.809 0.711 0.572 0.634 0.700 0.690 0.695
Dynamic 3-shot + Automatic Prompt Optimization 0.829 0.821 0.825 - - - - - -
Dynamic 3-shot + Instruction Retrieval from Annotation Guideline 0.849 0.841 0.845 0.789 0.567 0.659 0.803 0.734 0.767
Two-step Inference 0.829 0.872 0.850 0.761 0.622 0.684 0.771 0.744 0.757
Two-step Inference + Instruction Retrieval from Annotation Guideline 0.920 0.885 0.902 0.733 0.789 0.760 0.843 0.780 0.810
Baseline (SOTA Foundation model) 0.939 0.942 0.940 0.856 0.876 0.866 0.882 0.912 0.897
Table 5: Results on Relation Extraction Tasks
BC5CDR (RE) BioRED
Precision Recall F1 Precision Recall F1
Basic Instruction 0.639 0.679 0.618 0.323 0.259 0.274
Dynamic 3-shot 0.662 0.646 0.654 0.392 0.292 0.318
Dynamic 3-shot + Instruction Retrieval from Annotation Guideline 0.632 0.656 0.647 0.615 0.417 0.497
Two-step Inference 0.732 0.801 0.741 0.474 0.354 0.406
Two-step Inference + Instruction Retrieval from Annotation Guideline 0.761 0.833 0.781 0.613 0.528 0.567
Baseline BioBERT: 0.693 BioGPT: 0.495 0.615 0.432 0.650 0.462 0.571 0.605 0.589
The Dynamic 3-shot approach outperformed the basic instruction method (which used prompts similar to those in prior studies, supplemented with randomly selected 3-shot examples) across all datasets except HoC. This improvement can be attributed to that more contextually relevant examples provided the LLM with additional background knowledge and annotation style about similar queries. However, while Dynamic few-shot does not involve model training or parameter updating, it still leverages manually annotated datasets. Conversely, in the Two-step Inference (Reasoning Step & Structurization Step) approach, few-shot prompting was not utilized due to the impracticality of pre-writing reasoning process for every instance in the training set. Nevertheless, the Two-step Inference method achieved significant performance improvements across all tested datasets. This highlights again that the commonly used structuralization for LLMs in biomedical text mining tasks can severely constrain the reasoning capabilities that frontier LLMs acquired during their training phase.
We also experimented with automatic prompt optimization, iteratively refining prompts by using natural language to describe the distributional differences between predicted and correct labels (resembling a "pseudo-gradient descent" approach). However, the performance gains were modest compared to the huge computational cost, likely because the discrepancy patterns between predicted and correct labels varied across test instances. As a result, a short natural language "gradient" struggled to generalize across the test set. Analyzing the prompts generated during optimization revealed that many included generic but dataset-irrelevant instructions (e.g., rewriting prompts in Markdown, adding overly detailed background introductions) or directly retailed the pseudo-gradient information, such as "the predicted label distribution on the test set was …, while the actual distribution was… Use this information to …" or "Example discrepancy in the last inference:…" Such instructions were clearly unusable by the LLM during inference. Nevertheless, some iterations produced potentially helpful guidance, such as "Be mindful of over-predicting topics such as …. These should only be selected if …". Given the computational cost and limited performance improvement, we cannot recommend this approach for production scenario in biomedical text mining. However, it could serve as a source of inspiration when manually curating prompts.
Instruction Retrieval from Annotation Guidelines generally improved performance, except when combined with Dynamic 3-shot on the BC5CDR (RE) dataset. Its impact was particularly notable for datasets like BioRED, which involve complex entity-relation annotations, or when paired with the Two-step Inference approach. Unfortunately, we could not apply this method to datasets such as LitCovid and HoC due to the absence of detailed annotation guidelines. Across all datasets, combining Two-step Inference with Instruction Retrieval from Annotation Guidelines consistently enhanced GPT-4o’s performance. However, this combination required extracting information and reasoning separately for each entity class, increasing the API usage overhead.
Overall, implementing systematic enhancements targeted for biomedical text mining enabled state-of-the-art LLMs to outperform baseline models on two datasets with minimal reliance on manually annotated data and achieve competitive performance on others. This challenges the prevailing notion in prior studies that LLMs underperform in discriminative tasks within the biomedical domain.
4.2 Evaluation of Reasoning Model (o1-mini)
Table 6 presents our experimental results on the reasoning model (OpenAI o1-mini). Except for the fact that the model’s built-in test-time computation replaces the need for Two-Step Inference, all prompting settings remain the same as those used with GPT-4o. Across all datasets, o1-mini performs slightly worse than GPT-4o enhanced with our proposed approach. We speculate that larger models, such as o1 or Deepseek-R1[72], which are capable of deeper reasoning, could achieve higher performance than o1-mini. However, due to computational costs and API providers’ constraints on processing time, we were unable to include these models in our experiments.
Table 6: Comparison of o1-mini and GPT-4o on various datasets.
o1-mini GPT-4o
Dataset Precision Recall F1 Precision Recall F1
LitCovid 0.824 0.895 0.837 0.831 0.909 0.873
HoC 0.770 0.849 0.807 0.827 0.865 0.841
BC5CDR-Chemical 0.767 0.828 0.797 0.920 0.885 0.902
BC5CDR-Disease 0.681 0.683 0.708 0.733 0.789 0.760
NCBI-Disease 0.777 0.763 0.770 0.843 0.780 0.810
BC5CDR (RE) 0.685 0.722 0.703 0.761 0.833 0.871
BioRED 0.465 0.448 0.456 0.613 0.528 0.567
  • Since the reasoning model does not require "Two-Step Inference" to retain the thought process, we utilized Dynamic 3-shot prompting for multi-label text classification tasks (LitCovid, HoC), and "Instruction Retrieval from Annotation Guideline" prompting for NER tasks (BC5CDR-Chemical/Disease, NCBI-Disease) and RE tasks (BC5CDR RE, BioRED).
4.3 Result of Model Distillation
We followed the same procedures as outlined in the original studies of HoC and BC5CDR-RE datasets to collect papers published after April 2024, and then constructed synthetic training datasets entirely annotated by LLM. These datasets were used to fine-tune the BioLinkBERT-Large model, which was subsequently evaluated on a manually annotated gold-standard test set. As shown in Table 7 , the distilled BioLinkBERT model achieved a Macro-F1 score of 0.817 on the HoC dataset, slightly behind the performance of 0.873 when fine-tuned on the official training set which is smaller but annotated by human experts. For the BC5CDR-RE dataset, the distilled BioLinkBERT model achieved a Macro-F1 score of 0.646, closely approaching the performance of 0.650 obtained by fine-tuning on the official training set. Notably, it outperformed the fine-tuned BioGPT, which achieved a Macro-F1 score of 0.462 and was reported as the SOTA method on this dataset.
Table 7: Results about Model Distillation on HoC and BC5CDR-RE Datasets
HoC BC5CDR-RE
Precision Recall F1 Precision Recall F1
Distillated Model (BioLinkBERT-Large Fine-tuned on Augmented Dataset) 0.797 0.835 0.817 0.625 0.668 0.646
Baseline (BioLinkBERT-Large Fine-tuned on Trainset) 0.887 0.864 0.873 0.693 0.615 0.650
Moreover, we investigated the relationship between the size of supervised data and the performance of BERT-based model. As shown in Figure 6, the performance of BioLinkBERT improves rapidly with an initial increase in supervised data. However, the benefit of additional data diminishes beyond a certain level, and the performance approaches a plateau. Manually annotated datasets enable the BERT model to achieve higher performance with relatively smaller data volumes. Nonetheless, leveraging large-scale data labeled through stacked prompts from LLMs can yield performance levels close to those obtained with manual annotation, while significantly incurring lower human resource costs.
Based on these findings, we propose that in production scenarios requiring continuous inference of a large amount of texts but with less stringent performance demands—such as document classification to assist with reading—LLMs guided by targeted prompt engineering have the opportunity to replace human annotators in the widely adopted “manual annotation–model training–inference” workflow. This approach offers a cost-effective alternative while maintaining acceptable levels of performance for many practical applications.
Figure 6: Relationship between Training Data Size and Performance on HoC Dataset
5 Conclusions and Future Work
In our study, we first analyzed the factors contributing to the suboptimal performance of LLMs when applied to biomedical text mining tasks using conventional approaches. Building on this analysis, we developed a composite prompting strategy specifically targeted to address these challenges, and our results demonstrate that the frontier LLMs like GPT-4o could challenge the prevailing notion that they are inherently unsuitable for discriminative tasks in biomedical domain. While core ideas of these prompt engineering techniques and information retrieval frameworks were not originally ours, we customized them to better suit the specific needs of the field.
Our results indicate that combining two-step inference and instruction retrieval from annotation guidelines can greatly enhance LLM performance. However, this method relies heavily on predefined rules for segmenting the annotation guidelines and extracting relevant information during inference, demanding significant preparatory and coding efforts. Future improvements could include developing an LLM annotator agent that automatically queries the necessary guidelines and examples, or integrating a trainable retriever like in Li et al.’s research[73] to dynamically extract guidelines, instead of the rule-based approach. Additionally, we did not fine-tune or deploy any open-source LLM due to computational limitation. Nonetheless, the proposed prompt engineering strategies and model fine-tuning approaches are complementary, and their combination may potentially yield even better performance.
In production applications, for tasks requiring highly accurate data annotation—such as integrating information extracted from literature data into a biological database—LLMs, like existing machine learning methods, cannot guarantee performance at the level of human experts. Even though we anticipate that future iterations and upgrades of foundational models (including reasoning models) will further enhance LLM performance in biomedical text mining, reaching human expert-level accuracy remains a significant challenge. Nonetheless, the inference process of LLMs can serve as a reference for human annotators, potentially reducing their workload.
Furthermore, for applications that can tolerate some degree of annotation errors (e.g., improving literature readibility), LLMs, despite performance variations, still demonstrate practical predictive capabilities. One of the key advantages of LLMs is their ability to "cold-start" when new annotation needs arise—such as new annotation objects, types, or texts—without relying on supervised data constructed by human annotators. However, the inference cost of LLMs can become a concern when annotating large volumes of text, such as labeling all papers in the PubMed database, making model distillation a practical alternative.
Additionally, since our evaluations were conducted on commonly used biological text databases, we cannot entirely rule out the risk of data contamination (Appendix 7.3). Considering this, we recommend continuous monitoring and management of LLM performance when deploying this methodology in production environments.
6 Data Availability
The data and prompt used in this study can be found at the following repository: https://github.com/ekkkkki/LLM-Replace-Annotators-in-Biomedical-Text-Mining.